OpenStudy (asapbleh):

A movie star, unwilling to give his age, posed the following riddle to a gossip columnist: "Seven years ago, I was eleven times as old as my daughter. Now I am four times as old as she is. " How old is the movie star?

OpenStudy (anonymous):

If you take his present age as x and his daughters age as y x=4y (his present age is 4 times that of his daughters) and also x-7=11*(y-7) (from what he said about his age 7 years ago) solve these two equations you will get the answer

OpenStudy (asapbleh):

ok i will now try to solve

OpenStudy (asapbleh):

I have solved the equation you have gave me. I got y(daughter) as 12 years old present time, x(movie star) as 48 years old present time, this does not seem to be the correct answer because I have looked in the book and it is different?

OpenStudy (anonymous):

the answer should be 40

OpenStudy (asapbleh):

yes, how did you get 40 as your answer?

OpenStudy (anonymous):

just by solving the equation I gave

OpenStudy (asapbleh):

x=4y x-7=11y(y-7) 4y-7=11y-77 84=7y 12=y 48=x

OpenStudy (anonymous):

it wont be 84 = 7y it would be 70=7y

OpenStudy (asapbleh):

oyeah, lol how could of i got basic math wrong. thankyou for your help, it was very much appreciated. mind helping me with one more problem?

OpenStudy (anonymous):

you are welcome. Sorry I have to leave. Good luck with you question. Bye...

OpenStudy (asapbleh):

thankyou anyway